11- $2.56 Iterated Square Root Consider a sequence an. given recursively as follows: 2+ an-1: 2 2 Show that a,n 2 as n x
Answer #1

Solution- 94 = 2 and an = JZ ant 67/2 show that an - 2 98 n a Now 91= E 92 = 2+91 = 12+ Tz <2 9ca2 93= 52+02 = 2+52 +52 <2 a